### v1

# Biomedical Image Segmentation: A Systematic Literature Review of Deep Learning Based Object Detection Methods

Abstract
Biomedical image segmentation plays a vital role in diagnosis of diseases across various organs. Deep learning-based object detection methods are commonly used for such segmentation. There exists an extensive research in this topic. However, there is no standard review on this topic. Existing surveys often lack a standardized approach or focus on broader segmentation techniques. In this paper, we conducted a systematic literature review (SLR), collected and analysed 148 articles that explore deep learning object detection methods for biomedical image segmentation. We critically analyzed these methods, identified the key challenges, and discussed the future directions. From the selected articles we extracted the results including the deep learning models, targeted imaging modalities, targeted diseases, and the metrics for the analysis of the methods. The results have been presented in tabular and/or charted forms. The results are presented in three major categories including two stage detection models, one stage detection models and point-based detection models. Each article is individually analyzed along with its pros and cons. Finally, we discuss open challenges, potential benefits, and future research directions. This SLR aims to provide the research community with a quick yet deeper understanding of these segmentation models, ultimately facilitating the development of more powerful solutions for biomedical image analysis.
Keywords Biomedical Image Segmentation $\cdot$ Biomedical Imaging Modalities $\cdot$ Deep Learning $\cdot$ Object Detection based Segmentation $\cdot$ Systematic Literature Review
1 Introduction
Biomedical image segmentation plays a vital role in the analysis and diagnosis of diseases related to various organs. In this regard, the accurate identification of the region of interest in the biomedical images is a significant phase for identifying affected regions in the image. There are several alternative methods to achieve this segmentation. object detection based methods is a type of instance segmentation that is very common in this domain. Deep learning-based object detection methods offer significant advantages over the other alternatives. Considering the prominence of deep learning based object detection-based methods for biomedical image segmentation, we believe a comprehensive review of this field can be a a valuable work to provide a quick insight into the field. Although, there are few reviews (section 1.2) in the related topic, there is no standard review on the target direction.
Inspired by the systematic literature review (SLR) methodology (section 2) from software engineering domain [1], this paper presents a comprehensive SLR of the object detection based deep learning methods used for biomedical image segmentation. Following SLR guidelines, we collected 150 articles (primary studies) in an unbiased manner. After collecting the primary studies, we extracted the results. The results include the deep learning models, modalities, a list of diseases, and the metrics used for the analysis of the segmentation methods. We also provide critical analysis of the methods and presented each of the article with their pros and cons.
1.1 Background and Motivation
Segmentation of biomedical images can be categorized into two main categories including semantic segmentation and instance segmentation. In semantic segmentation, each pixel of an image is analyzed and classified into a particular class. This type of segmentation particularly helps in capturing the quantitative shapes of objects (in the case of biomedical imaging, a disease or affected region) in the image [2]. The deep learning models applied for semantic segmentation are typically found in three categories including region-based (RB) segmentation [3], fully convolutional neural network (FCN) [4, 5, 6], and weakly supervised or semi-supervised learning (SSL) approaches [7, 8].
In the instance segmentation, each pixel of image is assigned to some specific instance which has a significant role in biomedical imaging for the identification and analysis of various types of diseases or same type of disease with different locations in the image. Different types of algorithms are used for instance segmentation e.g., supervised learning (SL) models and semi-supervised learning (SSL) approaches. Each approach has its own drawbacks and benefits associated with them. Recently, fully supervised models have got proper attention for biomedical image instance segmentation which are based on human experts’ knowledge for annotations. These models are very laborious and expensive in terms of obtaining dense and accurate annotations  [9]. To overcome the problem of time consumption and human efforts associated with annotation for supervised learning approaches, some semi-supervised models have been proposed which require small set of annotated biomedical image samples as compared to the fully supervised instance segmentation models  [10]. The instance segmentation is mainly categorized into two main types namely proposal-free  [11] and proposal-based instance segmentation  [12]. The former category is based on morphological properties and instance-aware features whereas the latter is type is based on object detection approach which is further divided into two stage detectors (2SD), one stage detectors (1SD), and point-based detectors (PBD) or advanced detectors.
Biomedical image segmentation has gained a significant attention from researchers, resulting in the development of various methods across different categories. Among these, object detection-based methods utilizing deep learning have emerged as particularly effective for biomedical image segmentation. However, unlike primary studies, currently there is no comprehensive review of this specific topic (see subsection 1.2). To the best of our knowledge, there is no systematic literature review (SLR) on the topic. There are few review articles but they are not following some standard methodologies. Therefore, an SLR is essential to consolidate existing knowledge, identify research gaps, and provide a cohesive understanding of the advancements in this field. This review aims to bridge that gap by thoroughly examining and synthesizing the literature on deep learning-based object detection methods for biomedical image segmentation.
Table 1: COMPARISON WITH RELATED SURVEYS
S.No Authors Covered up to Year Imaging Modalities Processing Tasks Performed Models Targeted Diseases/Organs Covered
1 Isin, Ali et al.  [8] 2016 MRI (T1, T1-Gd, T2, FLAIR) Automatic, Semantic Segmentation CNN, ML, Generative Models Brain Tumor
2 Anwar, Syed Muhammad et al.  [13] 2018 CT, MRI, PET, X-Rays, Ultrasound Segmentation, Classification, Diagnosis CNN different variants + ML models Lungs, Breast, Brain, Thyroid
3 Haque and Neubert  [14] 2019 Clinical, X-Rays, CT, MRI, Ultrasound, OCT, Microscopic Images Semantic Segmentation, Classification, Analysis ML, CNN, RBM, Auto encoders, GANs, RNN, FCN, U-Net, ResNet, V-Net Brain, Breast, Liver, Spleen, Heart, Stomach, Ventricles, Arteries, Skin
4 Yahya and Boubakeur  [15] 2019 PET, CT, X-Ray, Ultrasound, Histological Images Classification, Segmentation VGG-16, GoogleNet, AlexNet, ResNet, ResUnet, SegNet, UNet++, RD-Unet, RNN, ML Models Not Mentioned
5 Tongxue Zhou et al.  [16] 2019 T1, T2, T1C, FLAIR, T1-DUAL, T2-SPIR, T1-IR, CT, PET, MRI Semantic Segmentation LeNet, ZFNet, GoogleNet, DenseNet, ResNet, AlexNet, U-Net, RNN, SVM, FCN, CRF Brain Tumor, Brain Stroke, Ischemic stroke lesion, Abdominal Organs, Intervertebral Disc
6 Kar et al.  [17] 2020 Mammograms, MRI, Endoscopic OCT, Microscopic images, Radiographs, Colonoscopy, CT, X-Rays, Fundus Camera, Multi-modal, 3D Images Semantic Segmentation U-Net++, Dual U-Net, SegNet, PSPNet, U-Net, ReNet, ResNet, VGG-16, AlexNet, GoogleNet, GAN, DBN, RBM, AE, SAE, RNN, FCN Breast, Brain, Eye, Pancreas, Gland, Chest, Head, Tissues, Lungs, Cells
7 Taghanaki, Saeid Asgari, et al.  [2] 2020 Microscopy, MRI, CT, EM, Fundus, X-Ray, PET Semantic Segmentation SegNet, FCN, Auto-DeepLab, RefineNet, ResNet-38, DeepLabV3 Prostate, Lung Nodule, Pancreas, Abdomen, Spleen, Breast, Heart, Spinal Cord, Retinal Vessel, Brain, Colon, Vertebral Body
8 Tajbakhsh, Nima, et al.  [18] 2020 MRI, CT, X-Rays, LGE, bSSFP, DRR, Ultrasound, Angiography, FLAIR, Fundus images, Histography Semantic Segmentation, Classification, Detection Self-Learning Models, Semi-Supervised Models, Unsupervised Models, Ensembled Models Lungs, Abdomen, Breast, Tumor, Brain, Pancreas, Kidney, Liver, Heart
9 Iqbal, Ahmed et al.  [19] 2021 CT, MRI, X-Ray, Mammography, Ultrasound Segmentation Vanilla-GAN, Cycle-GAN, Patch-GAN, Style-GAN, DC-GAN Retina, Brain, Skin, Heart, Breast, Bone, Lungs
10 Punn and Agarwal  [20] 2021 Ultrasound, PET, X-Ray, CT, FLAIR Segmentation U-Nets’ Variants (Better U-Nets, Inception U-Nets, Attention U-Nets, Ensemble U-Nets) Heart, Lungs, Chest, Bones, Blood Vessels, Soft Tissues, Skin
11 Imran et al.  [21] 2021-2022 CT, MRI, US, Microscopy, Histology, EM, PET, Colonoscopy Semantic Segmentation FCN, U-Net, Gamma-Net, SSIM, AFPNet, ResNet Brain, Chest, Breast, Skin, Prostate
12 Fernando and Tsokos  [22] 2022 3D-MRI Semantic Segmentation Statistical Methods, Deep Learning Models, Hybrid Models Brain
13 Xiao, Hanguang et al.  [23] 2023 MRI, CT Semantic Segmentation Different types of transformer models Abdomen, Heart, Brain, Lungs
14 Wang, R. et al.  [24] 2022 X-Rays, CT, MRI, Ultrasound Semantic Segmentation U-Net, 3D-Net, RNN, ResPath, FED-Net, FCN, CDNN, GAN, Transformer models Liver, Pancreas, Heart, Lungs, Brain, Kidney, Spleen, Skin, Thyroid, Prostate, Colon
15 Iqbal et al.  [25] 2023 CT, MRI, X-Rays, Echocardiograms Semantic Segmentation CNN, U-Net, DeepLabV3, ResNet-18, RF, NB, DT, LR, KNN, XGBoost Chest, Arteries, Heart
16 Ilesanmi  [26] 2023 Fundus Classification, Segmentation CNN, FCNN, VGG, U-Net, LeNet-5, M-Net Arteries, Veins, Retina
17 Our Survey (SLR) 2024 30 different modalities Object Detection Based segmentation Two-stage detectors, One-stage detectors, Advanced detectors Maximum Body Organs Covered
1.2 Related Surveys
Table 1 presents related surveys in the field and their comparison with our current SLR. Many surveys have been conducted related to different types of segmentation methods of biomedical images. All of them cover different types of biomedical image segmentation approaches with keeping in consideration various parameters in their scopes. As shown in Table 1, the major attractions of researchers in the last decade have been semantic segmentation using different types of machine learning models, semi-supervised learning-based models, and supervised learning-based models. A very less and insufficient attention has been paid to instance segmentation with further low attention towards object detection-based segmentation models.  Table 1 presents a comparison of all the existing surveys and distinction of our current survey. The main focus of our survey paper will be different types of object detection-based segmentation models applied for segmentation of biomedical imaging which will be the first survey conducted on this worth-attraction and important direction. Furthermore, it is evident that no proper attention has been given to the segmentation of biomedical imaging based on object detection models which are very powerful and recent techniques for the diagnosis of medical images.
Moreover, the existing surveys does not follow SLR or other standard methodology. SLR provides a structured predefined set of rules and comprehensive synthesis of research findings. It offers critical insights on a given topi, and thereby yields a well structured survey. It is unbiased as every step from searching to the final results presentation are clearly defined in an organized way.
1.3 Contributions
The key contributions include:
  • We conducted the first SLR on this topic, where we selected 148 primary studies as our final list and conducted a comprehensive analysis.
  • Various types of biomedical imaging modalities extracted from the literature are presented along with their pros and cons for their usage in the disease detection and diagnosis.
  • Deep learning models based on object detection algorithms used for biomedical image segmentation are extracted from the selected articles.
  • We identified various challenges of object detection-based segmentation methods for biomedical images, and listed few directions for future research with a sketch of potential solutions.
2 Research Methodology
Inspired by its significance, and quality, we followed the SLR guidelines [1] to conduct this review. SLR provides a predefined rules and thereby provides an unbiased framework for conducting a high-quality literature review. This methodology formulates the whole architecture of the review, providing step by step protocols. Our research methodology is further described in the following subsections.
2.1 Research Questions
Research questions are the key fillers of the research providing a baseline for analysis the quality and content of the research. This SLR is based on the following research questions (all in the domain of deep learning for objected detection biomedical image segmentation). Each question yields a list of outcomes with critical analysis among them.
  1. RQ 1:
    What are the different imaging modalities used for detection and diagnosis of human diseases?
  2. RQ 2:
    Which object-detection based deep learning models are applied for the segmentation of different biomedical imaging modalities?
  3. RQ 3:
    What are the major problems and issues associated with each object detection based method when applied for biomedical image segmentation?
  4. RQ 4:
    What types of human diseases have been detected in biomedical imaging using different segmentation models?
  5. RQ 5:
    What are the statistical metrics used evaluation of the segmentation methods, and models used in the domain?
2.2 Search Strategy
The search was carried out with the following search string.
Figure 1: Research Methodology, and Search Results.
(Biomedical OR Medical) AND Segmentation AND (“Deep Learning” OR “Machine Learning”) AND (“Object Detection”) AND (Disease OR Disorder OR Abnormality)
We searched the articles using the mentioned serach string in the following libraries:
  1. 1.
    ACM digital library acm.org,
  2. 2.
    IEEE Xplore ieee.org,
  3. 3.
    SpringerLink link.springer.com,
  4. 4.
    ScienceDirect sciencedirect.com, and
  5. 5.
In addition, we identified some article via manual search from other venues, such as Eurographics digital library diglib.eg.org/ and  arxiv.org.
All these databases have a vast diversity in the number of publications both from scientific field as well as clinical field. The first search of relative literature was performed in November 2023.
2.3 Article’s Selection Criteria
The articles were selected based in two stages. The primary selection was based on the title, abstract and the related keywords. All the articles were reviewed by authors for finding the relevance of inclusion or exclusion according to criteria specified. In this stage, we selected 340 research articles. The final selection is based on the reading the full paper. In this stage a total of 139 (out of 340) articles were selected. In addition, 24 articles were selected through snowballing and manual search, making a total of 163 articles, which were then passed through quality assessment in the next step. The articles were included/included on the following criteria.
  • We included peer-reviewed journal and conference papers, as well as non-peer-reviewed preprints, but excluded non-peer-reviewed reports and books.
  • We included the articles with major focus on object detection-based segmentation methods and excluded those related with general segmentation approaches and methods related specifically to clinical systems.
  • Previously conducted surveys and reviews were also excluded from the study.
  • Research articles related to approaches of semantic segmentation and methods based on proposal-free were also excluded.
  • Due to rapid development in the field of deep learning models and automatic segmentation methods in the last one decade, it is believed that including older articles will not be so fruitful and necessary. Resultantly, the research articles published after 2013 have been included and the older articles were excluded.
  • We excluded non-English articles.
  • We excluded articles on quality assessment.
2.4 Article’s Quality Assessment
In addition, to the above selection criteria, we excluded 17 articles based on their quality. To evaluate the quality and relevance of each article for our study, we used a quantitative metric indicating articles quality. The metric was calucalted for each article as:
  1. 1.
    Citation Frequency: Number of citations per year on Google Scholar.
    • Good (1): 8+ citations per year (or articles published after 2020)
    • Average (0.5): 5+ citations per year
    • Poor (0): Less than 5 citations per year
  2. 2.
    Publication Venue: Quality and relevance of the journal or conference.
    • Good (1): Published in a high-quality, relevant journal or conference
    • Average (0.5): Published in a multidisciplinary venue
    • Poor (0): Published in a low-quality or irrelevant venue
  3. 3.
    Methodology: Novelty and article presentations.
    • Good (1): Novel and well written/presented.
    • Average (0.5): Moderately novel and/or Moderately presented
    • Poor (0): Lower novelty and/or poor writing/presentation
  4. 4.
    Results Analysis: Accuracy and clarity of results analysis.
    • Good (1): Thorough and accurate analysis with clear comparisons or theoretical proofs
    • Average (0.5): Moderate analysis with some clarity issues
    • Poor (0): Inadequate analysis
Each question was rated with either 1, 0.5 or 0; and then the accumulative score was used as quality metric. There were 15 articles scoring the accumulative value below 2 and they were excluded from our selection.
2.5 Extracted Information from the Papers
Information strictly related to object detection based biomedical imaging from each paper was extracted which includes following specifications.
  • Any pre-processing method applied for enhancing the quality of biomedical images before actual processing by the relative models.
  • The overall methodology adopted for automatic segmentation of biomedical images.
  • The implementation details including programming languages, software tools, applications, and hardware specifications.
  • Dataset’s description including total size, training and testing ratios, and cross validation if available.
  • Performance evaluation metrics
  • Comparative analysis with other models if they have performed and available.
Different manuscripts have performed their performance evaluation using different parameters but few of these standard parameters include accuracy, True Positive Rate (TPR, also called sensitivity), True Negative Rate (TNR, also called specificity), False Discovery Rate (FDR), Jaccard Index (JI), Dice Similarity Coefficient (DSC), Hausdorff Distance (HD), Intersection Over Union (IOU), area under receiver operating characteristic curve, and Kappa. Except the HD, the higher value of the metric indicates better performance and lower value indicates worse performance. In case of HD, lower value represents better performance and higher value shows worse performance  [27].
2.6 Presentation of the Findings
The information collected from the selected articles are categories into three categories. First, the imaging modalities and the deep learning models identified in the literature are presented in section 3 and section 4, respectively. Then, section 5 presents all the segmentation methods in detail, discussing their pros and cons, and summarizes various findings in tabular form. Section 6 highlights various research challenges and research gaps for future research. Finally, the paper is concluded in section 7.
3 BIOMEDICAL IMAGING MODALITIES
There are many types of biomedical images’ modalities applied for detection, diagnosis, and treatment of various types of diseases. Depending on the capturing mechanisms, materials and equipment used, and the procedure followed for their diagnosis, these images have been categorized in different classes which are all covered in this section. Various types of biomedical imaging modalities are shown in Figure 2.
Figure 2: Biomedical Imaging Modalities
Medical Radiology Images
Although, medical radiology is a major branch of medical sciences that deals with detection, diagnosis, and treatment of various types of diseases using imaging technology, based on some standardized characterization of few imaging modalities, they have been classified under this category. The important medical radiology imaging modalities include X-rays, Computer Tomography (CT), Magnetic Resonance Imaging (MRI), and angiography images which are special types of X-rays or CT images  [28].
Ultrasound Images (US)
The second important category of medical images modalities is ultrasound images which are used for diagnosis of various organs e.g., abdomen, kidneys, breast, spleen, liver, pancreas, and lungs. The important modalities of ultrasound are sonography images, color doppler, automatic breast ultrasound (ABUS), contrast enhanced ultrasound (CEUS), tissue elasticity, shear wave elastography (SWE), 3D ultrasound, and power doppler  [29].
paragraphMEndoscopy and Colonoscopy images The colonoscopy and endoscopy images are used for the diagnosis and treatment of gastrointestinal tract which starts from mouth, goes through pharynx and oesophagus, stomach and small intestine, large intestine and colon. The important categories of these images include white light endoscopy (WLE), narrowband imaging (NBI), optical coherence tomography (OCT), confocal laser endomicroscopy (CLE), endoscopic ultrasound (EUS), photoacoustic endoscopy (PAE), and photoacoustic endoscope ultrasound miniprobe (PAE-UE)  [30].
Ophthalmology Imaging Modalities
Ophthalmology is a branch of medical sciences that deals with the analysis, diagnosis, treatment and recovery of disorders associated with eyes. For the diagnosis and treatments of different eye parts, various types of imaging modalities are used which have different properties depending on what parts of eyes are considered and which types of diseases are targeted. The important imaging modalities used for the diagnosis and treatment of eyes include fundus photography, scanning laser ophthalmoscope (SLO), optical coherence tomography (OCT), and PAOM  [31].
Dermatology Imaging
Dermatology is a branch of medical sciences that deals with the skin. The field has responsibility of both the medicinal purpose as well as surgical purpose. There are many imaging modalities that used for diagnosis and treatment of skin related diseases e.g., dermoscopy (related to skin), trichoscopy (related to scalp and hair), and onychoscopy (related to nails)  [32].
Functional Imaging
MRI, CT, US, positron emission tomography (PET), and single photon emission computed tomography (SPECT) are functional images because these are used for analysing brain functionalities  [33].
Microscopic Imaging
The basic concept of microscopic imaging is that it uses microscope for capturing the images of different body parts at very deep and detailed cellular levels for the detection and diagnosis of various abnormalities in the organs. There are many types of microscopic imaging modalities but important of them include selective plane illumination microscopy (SPIM), structured structure illumination microscopy (SSIM), two photon intravital microscopy (2P-IVM), two photon micro endoscopy (2P-ME)  [34].
Histopathology Imaging
Histopathology is the study of the signs of the disease using the microscopic examination of a biopsy or surgical specimen that is processed and fixed onto glass slides. The images used in this process are called histopathology images. To visualize different components of the tissue under a microscope, the sections are dyed with one or more stains. The histopathology images are used for the diagnosis of different human orgnas e.g., breast  [35].
Molecular and Cellular Imaging
Molecular and cellular images are very small types of images which are used for studying the finer details of structures of molecules and cells. These images are actually the microscopic images but used for the study of cells and molecules e.g., red blood celles (RBC), whie blood celles (WBC), yeast, crystal, and epithelium  [36].
4 DEEP LEARNING MODELS
The applications of deep learning in biomedical images and videos analysis have been the focus of research community due to their capability of processing complex data in an efficient manner  [37]. The biomedical image processing tasks by applying deep learning models are divided into four categories namely image classification, object detection, instance segmentation, and semantic segmentation. In semantic segmentation, each pixel inside the image is assigned to a specific class e.g., each pixel in liver will be assigned to tumour, parenchyma, or blood vessels. In case of instance segmentation, each individual object of a specific class is separately identified  [38]. The object detection architectures are divided into three categories namely two-stage detection models  [39], single-stage detection models  [40], and points-based detection models (which are also known as advanced detectors by many researchers).
Figure 3: Object detection based deep learning models for biomedical imaging segmentation  [40]
The important two-stage detection approach has different models e.g., R-CNN, Fast-RCNN, Faster-RCNN, Mask-RCNN, Cascade-RCNN, FCN, and R-FCN. The one-stage well-known models are YOLO, SSD, YOLOv2/v3/v4, RefineDet, and RetinaNet. The points-based models include CornerNet, CornerNet-Lite, CenterNet, and FoveaBox as shown in Figure 3. All are discussed in detail in the coming sections along with their structure, function, applications in biomedical imaging, and their pros and cons  [40].
4.1 Two-stage Detection Models
As explained in previous sections, the two-stage detection algorithms follow two steps for detection of specific object in the images in common, and boundary boxed areas for possible diseases in specific relevance to this work. In this section, the two stage detection models are discussed in detail.
1) Region Based Convolutional Neural Network (R-CNN): R-CNN is a two-stage object detection model used for the identification of different objects in an image. In case of biomedical imaging, it is used for the identification of different normal and diseased parts in the image. The R-CNN has been introduced to decrease the number of regions for reducing computation task. The detailed architecture and working mechanism of R-CNN are shown in  [41].
2) Faster R-CNN: Faster R-CNN is one of the most widely applicable two stage detection approach consisting of few important components. It has data layer, standard CNN, region proposal network (RPN) layer, Region of interest (ROI), ROI pooling layer, fully connected layer, and the classification layer. In the data layer, the input images are fed to the network for processing. The structure and function of faster R-CNN have been explained in  [42]. 3) Mask R-CNN: The mask R-CNN is a deep learning model used for object detection-based segmentation of images in computer vision. The output of mask R-CNN comes in three forms namely the class label of the object, the boundary box of the object, and the mask. The standard architecture of mask R-CNN consists of mainly three components which are the network backbone, region proposal network (RPN), and the ROIAlign. The tasks of each of these modules are outlined in  [43].
4) Cascade R-CNN: As highlighted by the authors in  [44], cascade RCNN has been developed to resolve the issues of overfiting and intersection over union threshold value problem by training the detectors in a sequential manner with the increasing value of threshold in which the output of one detector is given as training to the next detector. This sequential training of detectors resolves both the issues of overfitting and the quality mismatch problem  [44].
5) Feature Pyramid Network (FPN): Detection of objects with varying scales is very difficult and challenging task especially for small objects. Pyramids of same image of different scales can be used for object detection but this is a very time-consuming task and requires a lot of memory. As an alternative, we can use pyramid of features of the image and this pyramid can be used for object detection. FPN has been created as a feature extractor for achieving high accuracy and speed of processing as explained in  [45].
6) Region Based Fully Convolutional Network (R-FCN): R-FCN is a two-stage object detection model consisting of region-proposals and classification of regions. In R-FCN, the regions are extracted by region proposal network which fully connected convolutional neural network. In the working procedure of R-FCN, the image is read by the convolutional layer which create feature maps. The feature maps are used for creation of region proposal network which consists of many regions of interest. The convolutional layers after the feature maps creates each region of interest. After the extraction of region of interest, pooling is applied to generate the vote for each region of interest  [46].
4.2 One-stage Detection Models
In contrast to two-stage detection algorithms which consists of two stages namely region proposals and classification or object detection, the one stage detection algorithm carries out the detection task in only one stage. It uses the concept of anchors and grid box for the localization of the objects inside the image. There are many one-stage algorithms e.g., YOLO, SSD, RefineNet, and RetinaNet. All these models are discussed in this section with finer detail.
1) You Only Look Once (YOLO): In the two-stage object detection algorithms e.g., Fast R-CNN, the model first identifies the regions of interest with the help of RPN and in the second stage each region of interest is given class labels for detection. In case of YOLO, both these functions are performed simultaneously with a fully connected convolutional layer. After YOLO was introduced in 2015, so far, few versions have been developed each having different pros and cons. The current variants of YOLO include YOLOV2 to YOLOV8. The working mechanism of YOLO consists of few stages as shown in  [47].
2) Single Shot Detector (SSD): SSD is a one-stage object detection model consisting of two parts namely the backbone and the SSD head. The backbone is a type of pre-trained convolutional neural network used for feature extraction. The SSD head is one convolutional layer of few convolutional layers that generates the bounding box and class of the object. The complete working procedure of SSD is shown in  [48].
3) RefineDet: It is a single stage detection model that carries out the object detection process in only one stage. The simple structure of RefineDet consists of two parts namely anchor refinement module (ARM) and object detection module (ODM). There are pre-defined anchor boxes with fixed sizes, ratios and locations. The ARM tries to remove negative anchors for reducing search space to classify and adjust the sizes and locations of anchor boxes for better initialization for the regressor  [49].
4) RetinaNet: It is another one-stage detection model consisting of one backbone and two subnetworks. The backbone is a convolutional neural network like ResNet used for feature map generation of the whole image. The first subnetwork performs the classification based on the outputs of backbone model whereas the second subnetwork is responsible for identification of bounding boxes and the class of each bounding box  [50].
Figure 4: A chart showing the models (light blue color) and modalities (light green color). The numeric values indicates the number of papers who have used these models/modalities.
4.3 Point Based Detection Models
There are some object detection algorithms which can applied in procedures sightly different both from one-stage detection techniques as well as two-stage detection techniques. All these algorithms have been placed in a separate category. Their working mechanisms and power are higher than both the already discussed two-stage detectors as well as one-stage detectors. Therefore, they have been categorized as point based detectors because they use specific points for detection of objects in inages. Some of these detectors are discussed in this section
2) CornerNet: In cornerNet, the bounding boxes of objects inside image are detected using pair key points namely top-left corner and bottom right corner. This process of finding keypoints is performed by using convolutional neural network. By using this procedure, the need for designing anchor boxes which are applied in previously discussed one-stage detecting approaches is eliminated  [51].
2) CornerNet-Lite: CornerNet-Lite is a combined architecture of two different variants of standard cornerNet namely CornerNet-Saccade and CornerNet-Squeeze  [52]. The first variant applies the concept of attention mechanism for eliminating the need of processing all pixels of an image whereas the main responsibility of the CornerNet-Squeeze is to introduce a powerful backbone architecture.
3) CenterNet: In this architecture, the objects are identified by finding their central points. From these central points, the other properties of objects like orientation, location, size, and even pose of objects are found. The resulting architecture also known as CenterNet is faster, simpler, and more accurate than the traditional bounding box-based object detection models  [53].
Foveabox: FoveaBox is flexible, accurate, and anchor free architecture for object detection. All the previous anchor-based object detection models utilize anchors to detect objects inside an image, their performance and ability of generalization are limited to the design of anchors. In contrast, the Foveabox directly identifies boundary boxes and the existence of objects in the image without refereeing to the pre-defined anchors  [54].
5 SELECTED ARTICLES AND THEIR ANALYSIS
In this section, we discuss all the selected articles with in-depth analysis of the extracted information from each article. Each article is briefly described with its pros and cons. The articles are classified in three categories including two-stage detection, one-stage detection, and point based. In addition, modalities and diseases concerned with each model are discussed in detail. The challenges faced by these models, and their pros and cons have also been presented. Figure 4 shows different models and concerned modalities in a chart; where as Figure 5 present a chart of the modalities and concerned diseases. Figure 6 shows the year-wise distribution of the selected articles.
5.1 Articles related to Two-stage Detection
As described in section 4.1, two-stage object detection models include R-CNN, Fast R-CNN, Faster R-CNN, Mask R-CNN, Cascade R-CNN, FPN, and R-FCN. This section discusses the application of these models for the segmentation of biomedical images.
For example, a modified CNN called region-based convolutional neural network (R-CNN) multi-task model is used for the segmentation of histological images to detect prostate cancer [55], and for the detection of subretinal hemorrhage in OCT images  [56]. MITOS-RCNN [57] is another approach that uses R-CNN for the detection of breast cancer in the histopathological images. The R-CNN has also been used for segmenting pathological images to detect various tumors such as Glioblastoma Multiforme (GBM), Head and Neck Squamous Cell Carcinoma (HNSCC), Non-Small Cell Lung Cancer (NSCLC), and Lower Grade Glioma (LGG) [58]. Similarly, mask scoring R-CNN [59] is used for breast cancer detection in automatic breast ultrasound (ABUS) imaging. Zhang et al.  [60] proposed an end to end cell R-CNN for the analysis of pathological images. Their segmentation is based on global convolutional neural network (GCN).
Two-stage CNN model [61] is used for the segmentation of hard exudate lesions in colour fundus images. Similarly, a combination of convolutional neural network and the level sets was proposed by Huang et al  [62] is used for the segmentation of skin images for the detection of lesions. Yuan and Lo  [63] proposed an enhanced version of convolutional neural network for the segmentation of dermoscopic images to detect lesions in the skin. The faster R-CNN has been used in another article [64] for the detection of cells in the microscopic images. Another alternative enhanced model of faster R-CNN [65] has been used for segmentation of angiography images to detect pulmonary embolism. Similarly, Huang-Nan et al  [66] used fast R-CNN in combination with transfer learning for foot images to detect diabetic foot ulcers. Faster R-CNN has also been used for the detection of polyp in the colonoscopy images [67], the segmentation of ultrasound images to detect thyroid nodules [68], segmentation of breast cancer [69], , and segmentation of oncologic FDG PET images to detect lesion  [70].
[…]
Figure 5: A chart showing the modalities (light green color) and diseases (yellow color). The numeric values indicates the number of papers with each modality.
The authors in  [79] applied mask R-CNN for the segmentation of ultrasound images. They introduced some modification in the RPN for improving the segmentation accuracy of the mask R-CNN model. Similarly, a combination of mask R-CNN and 3D U-Net was applied by the researchers in  [80] for the segmentation of CT images of pancreas. The authors in  [81] used mask R-CNN for segmentation of lungs nodule. The proposed model consists of backbone network, RPN, ROI Align and fully convolutional neural network. The authors in  [82] proposed a combination of mask R-CNN and holistically nested edge detection (HED) model for the segmentation of MRI and CT images of liver.
Similarly, mask R-CNN was used by Linqin et al for detection and segmentation to be used applied for 3D visualization of pulmonary nodule  [83]. Zhiyong et al.  [84] suggested a hybrid model of mask R-CNN and inception-V3 for the segmentation of ultrasound images to detect prostate cancer. In the proposed model sobel convolution has been applied for sharpening the input images. Yuanhao et al.  [85] performed the classification and segmentation of ultrasound images for detection of breast tumour. They applied superpixel elastic deformation approach for data augmentation and mask R-CNN for segmentation of the ultrasound images for possible identification of breast tumour.
Similarly, Jui-ying et al.  [86] proposed mask R-CNN for the segmentation of sonogram images to detect breast cancer. The architecture they developed consists of CNN for feature map generation, RPN for extracting regions of interest, ROI align for extracting features of each ROI. In the proposed model, the RPN has been combined with fast R-CNN. For the assessment of fetal growth in ultrasound images, Moccia et al.  [87] suggested an advanced version of mask R-CNN known as mask R2-CNN. Meng et al.  [88] suggested mask R-CNN for the segmentation and detection of adenomatous colorectal polyps in the colonoscopy images. Similarly, Jin et al.  [89] used mask R-CNN for the detection of early gastric cancer in the endoscopy images. Qadir et al.  [90] proposed mask R-CNN for the detection and segmentation of polyp in the colonoscopy images. Jung et al  [91] applied mask R-CNN for the automatic segmentation of nuclei in histopathology images. Santos et al.  [92] applied the mask R-CNN for the segmentation of fundus images to detect lesions in diabetic retinopathy images.
Table 2: TWO STAGE DETECTORS FOR BIOMEDICAL IMAGING SEGMENTATION
ID-Ref Year Modality Model Disease $||$ Organ Evaluation Metrics Results
A001   [58] 2022 Pathological R-CNN Tumours F1-Score, Dice, Hausdorff 0.8216, 0.7088, 11.3141
A002  [55] 2018 Histology R-CNN Prostate Cancer IoU, Accuracy 79.56%, 89.40%
A003  [56] 2021 OCT R-CNN, Faster R-CNN Subretinal Haemorrhage Sensitivity, Specificity, Accuracy 90.36%, 85.98%, 93.98%
A004  [59] 2021 Ultrasound R-CNN Breast Cancer DSC, JI, Hausdorff, MSD, RMSD, CMD 85.0%, 75.2%, 1.65, 0.49, 0.76, 0.67
A005  [60] 2021 Histopathological Images R-CNN GBM, HNSCC, LGG, NSCLC F1-Score, Dice, Hausdorff 0.8216, 0.7088, 11.3141
A006  [57] 2018 Histopathology R-CNN Breast Cancer Recall 0.955
A007  [61] 2024 Fundus CNN Exudate Lesions AUPRC, DSC, IOU 0.893, 76.6%, 62.4%
A008  [62] 2024 Dermoscopic Images CNN Skin Lesions Jaccard Index 77.0%, 85.1%
A009  [63] 2017 Dermoscopic Images CNN Skin disease Jaccard Index 0.765
A010  [64] 2017 Microscopic Images Faster R-CNN RBC, WBC, Yeast, etc. Recall, Precision, F1-Score 0.9773, 0.7768, 0.8656
A011  [65] 2021 Angiograms Faster R-CNN Pulmonary Embolism AP, Sensitivity, Specificity 85.88%, 87.30%, 86.29%
A012  [66] 2022 JPG Format + CSV files Fast R-CNN Diabetic Foot Ulcers Accuracy 90%
A013  [68] 2023 Ultrasound Faster R-NN Thyroid Nodules AP 50%, AP 75%, AP 50:5:95% 97.4%, 81.3%, 68.7%
A014  [71] 2023 CT MRI Faster R-CNN U-Net Heart Disease Dice, Jaccard, Hausdorff, Running Time 0.864, 0.767, 29.918, 4.1 Sec
A015  [67] 2021 Colonoscopy Faster R-CNN Polyp detection Precision, Recall, F1-Score 0.943, 0.925, 0.934
A016  [72] 2020 Endoscopy Faster R-CNN Cascade R-CNN YOLOv3 Gastric Disease mAP 37.20, 52.44, 44.49
A017  [69] 2021 Histopathology Faster R-CNN Breast Cancer Precision, Sensitivity, IoU, F1 Score 1, 0.03, 0.55, 0.05
A018  [70] 2019 PET Faster R-CNN Lesion Detection Accuracy, Sensitivity, Specificity, PPV, NPV 97.2%, 95.4%, 98.1%, 96.8%, 97.0%
A019  [73] 2023 Histopathology Faster R-CNN Mitotic Detection AP, AP 50%, AP 75%, AP small, AP medium, AP large 25.35, 50.31, 25.45, 17.6, 16.17, -1
A020  [74] 2016 Microscopic Images Faster R-CNN Cancer Cells Precision, Recall, AP, AUC 0.990, 0.990, 0.906, 0.661
A021  [75] 2020 CT Mask R-CNN Heart, Lungs DICE, JI 97.6%, 98.1%, 95.1%, 97.8%
A022  [76] 2018 Oral Images Mask R-CNN Cold Sore, Canker, Sore Dice 0.774, 0.714
A023  [77] 2020 Endoscopic Images Mask R-CNN Gastric Cancer Image Sensitivity Lesion Sensitivity 0.76, 0.96
A024  [78] 2023 MRI Mask R-CNN Knee, Hip Dice 0.76, 0.75
A025  [79] 2018 Ultrasound Mask R-CNN Follicles Accuracy 96.5%
A026  [80] 2021 CT U-Net, Mask R-CNN Pancreas DSC, JI, REC, ACC 86.15%, 75.93%, 86.27%, 99.95%
A027  [81] 2018 CT Mask R-CNN Lungs mAP 0.7334
A028  [82] 2019 MRI CT Mask R-CNN Liver Dice CT: 0.94 MRI: 0.91
A029  [83] 2020 CT Mask R-CNN Chest Sensitivity, FP 88.7%, 88.1%
A030  [84] 2021 Ultrasound Inception V3 + Mask R-CNN Prostate Cancer mAP, DICE, IOU, AP 0.88, 0.87, 0.79, 0.92
A031  [85] 2019 Ultrasound Mask R-CNN Breast Cancer DICE, Accuracy, TP, TN 83.93%, 80.42%, 63.64%, 87.88%
A032  [86] 2019 Sonogram Mask R-CNN Breast Cancer mAP 0.75
A033  [87] 2021 Ultrasound Mask R-CNN Fetal growth AD, DI, DSC, H 1.95, -0.31, 97.90%, 1.5
A034  [88] 2020 Colonoscopy Mask R-CNN Adenomatous colorectal polyps mAP 50%, mAP 70%, mAP 75% 89.50%, 78.40%, 73.50%
A035  [89] 2022 Endoscopy Mask R-CNN Gastric Cancer Accuracy, Sensitivity, Specificity, PPV, NPV 90.25%, 91.06%, 89.01%, 92.61%, 86.81%
A036  [90] 2019 Colonoscopy Mask R-CNN Polyp Precision, Recall, Jaccard, Dice 94.1%, 86.1%, 73.2, 80.19
A037  [91] 2019 Histopathology Mask R-CNN Nuclei Segmentation F1 Score, Precision, Recall, Dice 0.913, 0.907, 0.923, 0.835
A038  [92] 2023 Fundus Mask R-CNN 4 Different Lesions AP, mAP 0.2515, 0.1548, 0.1042, 0.1577, 0.2687, 0.1589, 0.1274, 0.1388
A039  [93] 2020 Dermoscopic Images Mask R-CNN Skin Disease Precision, Recall, Average Precision 0.9021, 0.9187, 0.9085
A040  [94] 2021 Dermoscopic Images Mask R-CNN RetinaDeepLab Skin Disease Jaccard 80.04%
A041  [95] 2019 PET Mask R-CNN Lungs Tumour Precision, Recall, F1-Score 0.90, 1, 0.95
Table 2 (Continued) ID-Ref Year Modality Model Disease $||$ Organ Evaluation Metrics Results A042  [96] 2021 PET CT Mask R-CNN Tumour Prostrate ACD, VD, DSC 0.83+-0.91, -0.01+-+-0.79, 0.84+-0.09 A043  [97] 2024 Microscopic Images Mask R-CNN Cell Detection Dice, Accuracy, Precision, Sensitivity, F1 Score, Specificity, Volume Difference 0.862, 0.945, 0.901, 0.827, 0.862, 0.977, 0.082 A044  [98] 2023 Microscopic Images Mask R-CNN Anthrax Tissue Disease AP, AP75, IoU, DICE 0.560, 0.613, 0.751, 0.840 A045  [99] 2021 Microscopic Images Mask R-CNN Cell Segmentation SEG, DET, TRA, D-T 85.26%, 97.14%, 97.05%, 0.09 A046  [100] 2023 Ultrasound Cascade R-CNN, Faster R-CNN Arteries Disease Accuracy, AO, NRA, RASIRA 86.5%, 90.4%, 84.7, 88.8% A047  [101] 2022 Ultrasound Cascade R-CNN Thyroid Nodules mAP, TPR, FPR, TNR, FNR, Accuracy 87.1%, 0.9841, 0.0082, 0.9894, 0.0158, 0.9867 A048  [102] 2022 Ultrasound Cascade R-CNN Breast Cancer Accuracy, Jaccard, Precision, Recall, Specificity 97.51%, 72.47%, 82.70%, 99.13%, 79.35% A049  [103] 2019 Endoscopy Cascade R-CNN FPN Multi-class Artifacts Score Detection, Score Segmentation 0.3429, 0.3500 A050  [104] 2019 Endoscopy Cascade R-CNN Artifacts mAP, IoU, Score 0.3235, 0.4172, 0.3610 A051  [105] 2023 Ultrasound FPN Fetal head segmentation Heart DICE, IoU, Hausdorff, Conformity 89.08%, 80.70%, 19.76, 74.64% A052  [106] 2023 Endoscopy FPN Polyp Different Varying different varying A053  [107] 2021 Endoscopy FPN + CNN Polyp Average Precision, Segmentation Score 0.8986, 0.7771 A054  [108] 2015 Histopathology FCNN Breast Regions Average Precision, Average Recall, Average F1 Score, Variance Precision, Variance Recall, Variance F1 Score 0.91, 0.82, 0.85, 0.015, 0.02, 0.01 A055  [109] 2019 Fundus FCNN Diabetic Retinopathy DSC, Sensitivity, Specificity, Accuracy 96.7%, 98.1%, 99.9%, 99.9% A056  [110] 2018 Fundus FCNN Vessels Detection Accuracy, Sensitivity, Specificity 0.9210, 0.7215, 0.9576 A057  [111] 2017 Fundus FCNN Retinal Segmentation Sensitivity, Specificity, PPV, F1 Score 93.12%, 99.56%, 89.90%, 90.93% A058  [112] 2018 Biopsy FCN Dermis Epidermis RCPC, Jaccard 0.8801, 0.8845, 0.8900, 0.9962 A059  [113] 2018 PET CT FCN Tumour DSC, Classification Error, Volume Error 0.85, 0.33, 0.15 A060  [114] 2024 PET FCN Lymphoma Detection Sensibility, Dice 0.9953, 0.8665 A061  [115] 2020 Dermoscopy Colonoscopy Endoscopy U-Net Multiple mIoU, DSC, Recall, Precision 0.6255, 0.7649, 0.7156, 0.8007 A062  [116] 2023 Sonograms D-CNN U-Net Breast Cancer Accuracy, Sensitivity, Specificity, AUC, UIoU, MDice 97%, 97.7%, 96.4%, 0.96, 0.89, 0.92 A063  [117] 2019 Angiographic CT MRI CE-Net Lungs, Vessels, Eyes Sensitivity, Accuracy, AUC 0.8309, 0.9545, 0.9779 A064  [118] 2021 Colonoscopy ColonSegNet Polyp AP, IoU, DSC, Jaccard, Precision, Recall 0.8000, 0.8100, 0.7239, 0.8435, 0.8496, 0.9493
Huang et al.  [93] applied mask R-CNN for the segmentation of dermatology images to detect lesions in the skin. Similarly, Bagheri et al.  [94] suggested the application of mask R-CNN, Retina-Deeplab, and graph-based model for the segmentation of dermoscopic images to detect lesions in the skin. Zhang et al.  [95] applied mask R-CNN for the segmentation of PET images to detection tumour in human lungs. Similarly, Wang et al.  [96] applied dual mask R-CNN for the segmentation of PET and CT images to detect tumour and prostrate. Wang et al.  [97] applied mask R-CNN for the analysis of microscopic images. Similarly, Hoorali et al.  [98] proposed an enhanced version of mask R-CNN called URCNN for the segmentation of microscopic images to detect anthrax and some other tissue diseases. Similarly, Chen et al.  [99] suggested the use of faster R-CNN for the segmentation of microscopic images to detect cancer cells. The authors have applied faster R-CNN and the circle scanning algorithm for improving the detection of cancer cells.
Wang et al.  [100] tested three two-stages model namely cascade R-CNN, faster R-CNN, and double head R-CNN on ultrasound images of arteries for their segmentation to detect various abnormalities. Zheng et al.  [101] introduced and improved version of cascade R-CNN for the segmentation of ultrasound images to detect thyroid nodules. Chen et al.  [102] proposed cascade R-CNN for the segmentation of breast ultrasound images to detect lesions. Zhang and Xie  [103] suggested a hybrid model of FPN and cascade R-CNN for the detection and segmentation of different multi-class artifacts in the endoscopy images. Yang and Cheng  [104] suggested the implementation of cascade R-CNN for the detection and segmentation of artefacts in the endoscopic images. In another work, Chen et al.  [105] proposed a modified version of feature pyramid network (FPN) called Fourier feature pyramid network (FFPN) for the segmentation of ultrasound images. Su et al.  [106] proposed an improved version of FPN called FedNet for the segmentation of polyp in the endoscopic images. Wuyang et al.  [107] suggested FPN with other CNN networks for the detection and segmentation of polyp in the endoscopy images. Su et al.  [108], suggested an improved version of CNN named fCNN standing for fast scanning deep convolutional neural network for the segmentation and identification of regions in histopathology images of breast. Xue et al.  [109] proposed a multi-task segmentation approach for the segmentation of retina for the detection of diabetic retinopathy. A fully connected convolutional neural network was proposed by Li et al.  [110] for the segmentation of retinal fundus images to detect vessels in the images. Similarly, the fully connected convolutional neural network was applied by Feng et al.  [111] for the segmentation of retinal images. Pal et al.  [112] applied deep convolutional neural network (FCN) for the segmentation of biopsy images to detect dermis, epidermis, and non-tissue regions. Xhao et al.  [113] proposed a fully connected convolutional neural network for the segmentation of PET and CT images to detect tumour.
Pang et al.  [114] suggested CNN for the segmentation of PET images based on multi-temporal approach to detect lymphoma. For improving the performance of standard U-Net architecture, the authors in  [115] developed an improved version of the standard U-Net called double U-Net in which they stacked two U-Nets on the top of each other. The model was evaluated with four publicly available dataset of different modalities e.g., dermoscopy, colonoscopy, and microscopy. Yang et al.  [116] suggested the use of attention U-Net and deep convolutional neural network for the segmentation and classification of sonography ultrasound images for the detection of breast cancer. A model called CE-Net standing for context encoder network was applied for the segmentation of 2D medical imaging in the study presented in  [117]. JHA et al.  [118] proposed ColonSegnet for the detection, localization, and segmentation of polyp in the colonoscopy images.
5.1.1 Analysis of Two Stage Detectors
Different imaging modalities that have been the target of two stage detectors include CT images, MRI images, X-rays, ultrasound images, endoscopic images, colonoscopic images, dermoscopic images, histopathology images, pathological images, microscopic images, fundus images, and PET images. In the same fashion, the most frequently applied two stage detectors include mask R-CNN, Faster R-CNN, FCNN, FPN, cascade R-CNN, and RetinaNet, respectively. The most powerful, simple, accurate, and easily implementable models of these architectures is the mask R-CNN. If we critically analyse, the mask R-CNN model provide better results than other two stage detection models. In some cases, the faster R-CNN also provides comparatively better results than other models. Similarly, some fluctuations can also be observed in the results of other two stage detection models. The major factors behind these variations in the results of all models include the size of datasets been processed, the imaging modality, the architecture of the detection model, the pre-processing and post processing mechanisms applied, and the hardware and software specifications used in the experimentation.
Figure 6: Articles distribution on the basis of year of publication.
5.2 Articles related to One-stage Detection
One-stage object detection models for the segmentation of biomedical imaging include, YOLO, SSD, YOLOv1/v2/v3/v4, RefineNet, and RetinaNet. In this section, these models applied for the segmentation of biomedical images segmentation are discussed in detail along with other considerations. For example, Kavitha and Palaniappan  [119] applied YOLO for the segmentation of brain MRI images to detect tumour. Similarly, Chang et al.  [120] suggested two deep learning models being YOLO and fully connected convolutional neural network for the segmentation and classification of heart MRI images. Suryawanshi and Patil  [121] applied YOLOv5 for the segmentation of MRI images to detect brain tumour. Similarly, Moreno et al.  [122] applied a combination of two convolutional neural networks for the segmentation of left ventricle in heart MRI images. In the same fashion, Zhong and Deng  [123] introduced an improved version of YOLOv3 for the segmentation of MRI images for the detection of herniated intervertebral discs. Unver and Ayan  [124] used YOLO and Grabcut model algorithm for the segmentation of dermoscopic images to detect lesions in the skin. Similarly, Lemay  [125] applied YOLOv3 for the segmentation of 2D and 3D CT images for localizing the kidneys in the images and diseases including Cystic kidneys, hypertrophied kidneys, and tumoral kidneys.
Pang et al.  [126] suggested an enhanced model of YOLOv3 called YOLOv3-arch for the segmentation of CT images to identify cholelithiasis and classify gallstones. Similarly, Wang et al.  [127] applied the YOLOv2 and Fourier Photography Microscopy (FPM) in combination for the detection of WBC. Afshari et al.  [128] used YOLO to localize and detect different normal active organs in the 3D PET imaging modality. Similarly, Nguyen et al.  [129] used Faster R-CNN and YOLO for the segmentation of chest X-Ray images to detect different abnormalities. In a similar attempt, Liu et al.  [130] proposed the YOLOv8 for the X-Ray angiography processing to segment the coronary artery. Wu et al.  [131] used YOLOv5 for multi-type lesions detection in angiography images. Xu et al.  [132] used scaled-YOLOv4 for the detection of pulmonary embolism in the CT angiograms. Similarly, Mortada et al.  [133] applied YOLOv7 for the segmentation of echocardiographic images to identify and segment the anatomical structures of left heart. Pkhrel et al.  [134] suggested the use of YOLOv8 along with the augmentation for the segmentation of coronary artery in the angiograms to detect diseases. In another study, Osama et al.  [135] used YOLOv8 model for the segmentation of angiography images to identify and detect coronary artery stenosis. Zhuang et al.  [136] used YOLOv3 for the segmentation of left ventricle in the echocardiography. Similarly, Zhang et al.  [137] suggested the use of and enhanced model YOLOV3 one-stage segmentation model for segmentation of ultrasound images to detect the location of thyroid nodules.
Lu et al.  [138] used YOLOX for the segmentation of ultrasound images to examine the fetus for congenital heart disease. And enhanced model of YOLOV5 called YOLO-CSE was proposed by Hussein et al.  [139] for the segmentation of muscle ultrasound to detect myositis. Xie et al.  [140] presented YOLOV4 in their work for the segmentation of ultrasound images to detect different organs. In the same manner, Ghose et al.  [141] suggested the use of an enhanced version of YOLOV5 to segment the colonoscopic images for the detection of polyp. Doniorjon et al.  [142] used YOLOV4-tiny for the segmentation of endoscopic images to detect polyp. Kirthika and Sargunam  [143] applied YOLOV4 for the segmentation of endoscopic images to detect multi-class artefact. Durak et al.  [144] implemented different types of deep learning models for the detection of polyp in the gastro endoscopic images. Al Zorgani et al.  [145] proposed YOLOV2 for the detection of breast cancer in histopathological breast images. Rong et al.  [35] applied YOLO for the segmentation of nucleus and characterization of tumour microenvironment. Ayaz et al.  [146] proposed different deep learning models including mask R-CNN, U-Net, CNN, and YOLO for the segmentation of nucleus in the histopathological images. Narotamo et al.  [147] applied YOLO for detection of nucleus and U-Net for the segmentation of nucleus in fluorescence microscopy. Nair et al.  [148] applied YOLOV4 for the nucleus detection in the histopathology images of breast.
Similarly, Bsl et al.  [149] suggested YOLO for the identification of region proposal network to diagnose cancer in the breast histopathology images. Similarly, Pal et al.  [150] applied YOLOV3 for detection of lesions in the retinal fundus images. Lalitha and Padyana  [151] in fundus images using YOLO-RF which is a hybrid model of YOLO and random forest. Santos et al.  [152] used YOLOV5 for lesion detection in fundus images. Saha et al.  [153] implemented CNN with YOLO for the detection of glaucoma in the color fundus images. Ramachandran et al.  [154] proposed the application of YOLOv3 for the localization of optic disc in retinopathy fundus images. Bagheri et al.  [155] applied YOLO for lesion detection in skin using dermoscopic images. Singh et al.  [156] suggested fuzzy logic in combination with object detection models for the detection of cancer in the dermoscopic images. Yin  [157] applied YOLOV3 network for the segmentation of PET/CT to detect tumour in Lymphoma patients. Shwetha et al.  [158] performed a two-stages approach for the detection of pus cell and epithelial cell. Similarly, Kong and Shen  [159] proposed an enhanced version of YOLO for the segmentation of microscopic images to detect microorganisms.
Puerta et al.  [160] applied a combined model of YOLO and U-Net for the segmentation of microscopic images to characterize the neuron. Cao et al.  [161] suggested SSD and other segmentation models for the detection of benign and malignant breast cancer in the ultrasound images. Similarly, the SSD was applied by Kulhare et al.  [162] for the detection of different lungs abnormalities in ultrasound images. Ansari  [163] proposed a modified form of single shot detector named MP-FSSD (Multi-scale Pyramidal Fusion Single Shot Multibox Detector Network) for the segmentation of endoscopic images frames or the endoscopy to detect polyp. Chen et al.  [164] proposed SSD for the detection of Polyp in the gastrointestinal endoscopic images using the segmentation. In a similar attempt, Xu et al.  [165] suggested DAD-SSD (Dynamic Attention Deconvolutional Single Shot Detector) for the classification of different types of polyps in the colonoscopic images. De Lucena et al.  [166] applied RetinaNet for the detection of skin different diseases namely melanoma, dysplastic nevus, and healthy skin. Rehman et al.  [167] suggested deep RetinaNet for the segmentation of dermoscopic images to identify the melanoma lesion. A hybrid model of RetinaNet and ResNet was applied for the detection of skin cancer in dermoscopic images by Ahmed et al.  [168]. Bagheri et al.  [169] proposed three different object detection models for the identification of skin lesions in the dermoscopic images.
The detailed descriptions of these models along with the types of imaging modalities, the types of diseases these models have been used for the detection, the performance evaluation metrices and quantitative measures are shown in Table 3.
Table 3: One Stage Detectors for Biomedical Imaging Segmentation
Ref Year Modality Model Disease/Organ Evaluation Metrics Results
A065  [119] 2023 MRI YOLO Tumour Specificity, Sensitivity, Accuracy Jaccard 0.987, 0.952, 0.943, 0.955
A066  [120] 2018 MRI YOLO Heart Disease Dice, Hausdorff, Sensitivity 0.9193, 10.452, 0.9085
A067  [121] 2024 MRI YOLOV5 Brain Tumour Loss, Accuracy 0.03, 0.9832
A068  [122] 2019 MRI YOLO+CNN Heart Disease PGC, Dice, APD 98.59+-4.28%, 0.93+-0.06, 0.72+-0.62mm
A069  [123] 2019 MRI YOLOV3 Herniated Intervertebral Disc Loss, Execution Time 0.75-0.15, 23.9ms, 24.5ms, 24.3ms
A070  [124] 2019 Dermoscopic Images YOLOV3 Skin Disease IoU, Accuracy, Sensitivity, Specificity, Jaccard, Dice 90, 92.99%, 83.63%, 94.02%, 79.54%, 88.13%
A071  [125] 2019 CT YOLOV3 Kidney Disease Dice, IoU 0.851, 0.759
A072  [126] 2019 CT YOLOV3 Cholelithiasis gallstones Average Accuracy 92.7% for Stone, 86.50 for Cholelithiasis
A073  [127] 2018 Microscopic Images YOLOV2 WBC Detection Precision Rate, Recall Rate 1.0, 0.97
A074  [128] 2018 PET YOLO Active Organs Precision, Recall, ABBLE, WLE, IoU 75-98%, 94-100%, 14mm, 24mm, 72%
A075  [129] 2023 X-Ray YOLOV5 Chest Disease AP@.5, AP@.95 0.616, 0.322
A076  [130] 2023 X-Ray YOLOV8 Coronary Artery F1 Score 0.441
A077  [131] 2023 Angiograms YOLOV5 Lesion Precision, Recall, mAP@0.1, mAP@0.5 0.64,0.68, 0.66, 0.49
A078  [132] 2023 CT Angiograms YOLOV4 Pulmonary Embolism Average Precision 83.04%
A079  [133] 2023 Angiograms YOLOV7 Heart 3 Parts DSC 92.63%,85.59%,87.57%
A080  [134] 2023 Angiograms YOLOV8 Heart Disease F1 Score,mAP 0.35,0.59
A081  [135] 2023 Angiograms YOLOV8 Coronary Artery Stenosis Precision,Recall,mAP 82%,58%,65%
A082  [136] 2021 Angiograms YOLOV3 Heart Disease Speed,Dice,MAD,HD 2.1-2.25 fps,93.57%, 2.57mm,6.68mm
A083  [137] 2021 Ultrasound YOLOV3 Thyroid Nodule Mean Precision,Mean Recall 94.53%,95.00%
A084  [138] 2022 Ultrasound YOLOX Fetus Analysis AP,SD 0.835,0.146
A085  [139] 2023 Ultrasound YOLOV5+YOLO-CSE Mytosis Accuracy,AP 98%,96%
A086  [140] 2022 Ultrasound YOLOV4 Organs Detection AP,mAP 95.2%,91.59%
A087  [141] 2023 Colonoscopy YOLOV5 Polyp Precision,Recall,F1-Score,mAP@0.5 97.26%, 96.15%, 95.67%, 99.08%, 89.44%
A088  [142] 2022 Endoscopy YOLOV4 Polyp Accuracy 89.9% Training, 85.5% Testing
A089  [143] 2021 Endoscopy YOLOV4 Artefact mAP,Speed 49.82%,76ms
A090  [144] 2021 Endoscopy YOLOV4 YOLOV3 SSD Polyp mAP 86.39%,80.98%,55.06%
A091  [145] 2022 Histopathology YOLOV2 Breast Cancer F1 Score 0.839
A092  [35] 2023 Microscopic Images YOLO Nuclei Tumour Accuracy,Precision,Recall,F1 Score mIoU 0.7110,0.8308,0.6743,0.7409 0.8423
A093  [146] 2021 Histopathology YOLO Nucleus Segmentation F1 Score,mAP,mAR 0.954,0.954,0.954
A094  [147] 2019 Microscopic Images YOLO Nucleus F1 Score,Training Time 0.98,1.6 Sec
A095  [148] 2021 Histopathology YOLOV3 Breast Cancer F Measure 0.73
A096  [149] 2021 Histopathology YOLO Breast Accuracy, Precision, Recall, Specificity, FPR, FNR 0.9573, 1, 0.9239, 0.0761, 0, 0
A097  [150] 2020 Fundus YOLOV3 Lesions Average Precision 83.3%
A098  [151] 2021 Fundus YOLO-RF Diabetic Retinopathy Accuracy, Precision, Recall 99.33%, 97.20%, 99.10%
A099  [152] 2022 Fundus YOLOV5 Lesion mAP, F1 Score 0.1540, 0.2521
A100  [153] 2023 Fundus YOLO + CNN Glaucoma F1 Score, Sensitivity, Specificity, AUC 97.4%, 97.3%, 97.5%, 99.3%
A101  [154] 2020 Fundus YOLOV3 Retinopathy Accuracy 100%
A102  [155] 2020 Dermoscopy YOLO Skin Lesion Sensitivity, Specificity, Jaccard, bDice, Accuracy 95.1%, 99.5%, 99.20%, 92.30%, 85.92%
A103  [156] 2023 Dermoscopy YOLO Cancer Accuracy, Sensitivity, Specificity, Dice, Jaccard 95.16%, 90.86%, 97.27%, 86.06%, 92.51%
A104  [157] 2021 PET YOLOV3 Tumour Mean Performance 35%
A105  [158] 2023 Microscopic Images YOLO Pus Cell Detection mAP, Accuracy 0.87, 94.5%
A106  [159] 2023 Microscopic Images YOLO Microorganism Detection Precision, Recall, mAP@0.5 92.3%, 92.8%, 93.7%
A107  [160] 2023 Microscopic Images YOLO + U-Net Neurons Characterization MACC, MLO, MIoU, SDACC, SDLO, SDIoU 0.9714, 0.0755, 0.8026, 0.0224, 0.0542, 0.1593
A108  [161] 2017 Ultrasound SSD Breast Cancer APR, ARR, F1 Score 97.20%, 70.56%, 81.76%
A109  [162] 2018 Ultrasound SSD Lungs Diseases Sensitivity, Specificity 93.6%, 96.5%
A110  [163] 2022 Endoscopy SSD Polyp mAP 92.67%
A111  [164] 2021 Endoscopy SSD Polyp mAP 95.74%
A112  [165] 2023 Colonoscopy SSD Polyp mAP, Accuracy 76.55%, 74.40%
A113  [166] 2020 Dermoscopy RetinaNet Melanoma Dysplastic Detection Accuracy 68.8%, 72.5%
A114  [167] 2022 Dermoscopy RetinaNet Melanoma Average Precision 97%
A115  [168] 2023 Dermoscopy RetinaNet + ResNet Skin Cancer Jaccard, Dice, Sensitivity, Specificity, AUC 91.4%, 90.70%, 91.60%, 97.80%, 97.60%
A116  [169] 2022 Dermoscopy RetinaNet Skin Lesion Sensitivity, Specificity, Jaccard, Dice, Accuracy 88.56%, 96.25%, 80.02%, 87.62%, 94.37%
5.2.1 Analysis of One Stage Detectors
Like two stage detection models, the modalities that have been targeted by one stage detection models include MRI, CT, PET, X-Rays, dermoscopic images, microscopic images, ultrasound images, endoscopic, colonoscopic, histopathology, fundus images, and few other modalities. Among the one stage detection models, YOLO and its different versions are the most powerful and frequently applied models. The most frequently YOLO version applied for medical image segmentation is YOLO version 3 followed by YOLO version 5, YOLO version 4, and YOLO version 8, respectively. Few authors have applied YOLO version 2 and YOLO version 7 as well whereas some authors have applied SSD, RetinaNet, and RefineNet for the segmentation of biomedical images but their frequency is very less. The quantitative measures of different performance evaluation parameters are shown in the last column of the table in which many variations can be observed. The variations in the results are due to different reasons including easy processing and implementation, less computational cost, high accuracy, simple architecture, and strength for handling complex datasets. All these factors lead YOLO and its versions to overperform other one stage models.
5.3 Articles Related to Point-Based Detection Models
There are some object detection models for image segmentation which have some different properties both from one-stage detection models and two-stage detection models. Keeping in consideration these properties, the models have been placed in a different category which has been named as advanced object detection models. The most common models included in this category include CornerNet, CornerNet-Lite, objects as points, and Foveabox. In literature, many authors have applied these object detection models for the segmentation of biomedical imaging. For example, Nawaz et al.  [170] applied CornerNet for the segmentation of brain MRI to analyse it for brain tumour detection. Similarly, Nawaz et al.  [171] applied a combined model of CornerNet and Fuzzy K-means for the segmentation of dermoscopic images to detect melanoma in the skin. Similarly, Wang et al.  [172] proposed the use of CornerNet for the segmentation of colonoscopic images to detect polyp. In another study, VASILEIOS ALEVIZOS  [173] proposed a backbone of CornerNet for the segmentation of brain MRI to detection brain tumour. ALADHAD and MAHUM  [174] proposed CornerNet of the extraction of informative features from the X-ray images to segment it for the detection of knee osteoarthritis.
Sandhya et al  [175] presented CornerNet for the segmentation of MRI images to detect two types of malignant cancers namely low-grade glioma and high-grade glioma. Choi et al.  [176] suggested two deep learning models namely centernet and U-net for detection and segmentation of endoscopic images to detect gastrointestinal disease. Zheng et al  [177] proposed an improved version of Centernet for the segmentation of microscopic images to detect white blood cells in them. In the same manner, Jiang et al.  [178] proposed a modified version of CenterNet for the segmentation of colonoscopic images to detect polyp. In another work, Wang et al.  [179] suggested CenterNet for the segmentation of Optical Coherence Tomography Angiography (OCTA) images to detect retinal vascular bifurcation and crossover points. Similarly, Masood et al  [180] applied CenterNet for the segmentation of brain MRI images to detect tumour.
Similarly, Albahli and Nazir  [181] suggested CenterNet for the segmentation of X-Ray images to classify and localize the diseases in them. In another study, Thangaval and Palanichamy  [182] applied CenterNet for the segmentation of CT images to detect pulmonary cancers. Jasitha and Pournami  [183] analysed both the faster R-CNN and CenterNet for the segmentation of histopathology images to detect Glomeruli. In another study, Nazir et al.  [184] applied CenterNet for the segmentation of retinal images to detect diabetic retinopathy. Shen et al.  [185] applied the combination of R-CNN, CenterNet, and deep snake for thyroid nodule detection in the ultrasound images. Mattikalli et al.  [186] developed an ensemble model of VFNet, FoveaBox, and RetinaNet for the detection of lesion in CT images. Similarly, Mathai et al.  [187] applied a hybrid model of VFNet, FoveaBox, FCOS, and DetectoRS for detection of lymph nodes in MRI. Sakunpaisanwari et al.  [188] different two stage detector models and one stage detector models including CornerNet and CenterNet for blood vessels detection in CT images. Durak et al.  [144] implemented different two stage detectors, one stage detectors, and other advanced detection models including CenterNet, for the detection of gastric polyps in the endoscopic images. Mathai et al.,  [189] applied an ensembled model of foveabox with other one stage detectors to detect lymphadenopathy in MRI images. Similarly, Celik et al.  [190] studied the role of different deep learning models including foveabox in the detection of lesions on panoramic radiographs. Cho et al.  [191] compared the performance of RetinaNet with cornernet and centernet in automatic tip detection of surgical instruments in biportal endoscopic spine surgery. In another work, Sakunpaisanwari et al.  [188] applied different deep learning models including centernet and cornernet for blood vessels detection in CT scans of lower extremities.
Nguyen et al.  [192] applied different deep learning models including centernet and cornernet for the detection of glomeruli and nuclei in pathological images. In another work, Zheng et al.  [177] applied centernet for the detection of blood cells in the microscopic images. In another work, Su et al.  [193] presented a centernet based on heatmap pyramid structure for the detection of rib fracture in CT images of chest. Kitvimornat et al.  [194] applied faster R-CNN, RetinaNet, and Centernet for the detection and characterization of parasite eggs in the microscopic images. Similarly, Jiang et al.  [195] suggested the use of centernet along with Resnet 50 for feature extraction in the diagnosis of ultrasound images. In another attempt, Dai et al.  [196] suggested the use of YOLOv4 and Centernet for the diagnosis of breast ultrasound to detect lesion. Wang et al.  [172] applied few deep learning models including centernet for the detection of polyp segmentation in colonoscopy. Similarly, Ma et al.  [197] proposed a transformer based centernet for the liver tumor detection in CT images. The detailed descriptions of all these models along with other specifications are shown in Table 4.
Table 4: Advanced detectors for biomedical imaging segmentation
ID-Ref Year Modality Model Disease/Organ Evaluation Metrics Results
A117  [170] 2021 MRI CornerNet Brain Tumour Accuracy, Precision, Recall 98.7%, 97.4%, 96.9%
A118  [171] 2022 Dermoscopic Images CornerNet Skin Disease mAP, Sensitivity, Specificity, Accuracy 99.48%, 99.39%, 99.63%, 98.8%
A119  [172] 2022 Colonoscopic Images CornerNet Polyp Precision, Recall 99.36%, 96.4%
A120  [173] 2023 MRI CornerNet Brain Tumour Accuracy 89.8%
A121  [174] 2023 X-Ray CenterNet Knee Osteoarthritis Accuracy 99.14%
A122  [175] 2022 MRI CornerNet Cancer Accuracy, Precision, Recall 91.3%, 91%, 88%
A123  [176] 2020 Endoscopic Images CenterNet, U-Net Gastrointestinal Disease mAP 0.1932
A124  [177] 2023 Microscopic Images CenterNet WBC Detection mAP 98.8%
A125  [178] 2023 Colonoscopic Images CenterNet Polyp Precision, Recall, F1-Score, Average Precision 87.7%, 84.2%, 85.8%, 82.2%
A126  [179] 2021 OCTA CenterNet Retinal Vascular Average Precision, Crossover Points, mAP 80.81%, 85.86%, 83.34%
A127  [180] 2022 MRI CenterNet Brain Tumour Precision, Recall, F1 Score, Accuracy 99.52%, 99.92%, 99.72%, 98.07%
A128  [181] 2022 X-Ray CenterNet Different Organs AUC, Average IoU 0.888, 0.801
A129  [182] 2024 CT CenterNet Pulmonary Cancer DSC, Sensitivity, PPV 99.07%, 99.65%, 98.96%
A130  [183] 2023 Histopathology Images CenterNet Glomeruli IoU, mAP 64.2%, 65.7%
A131  [184] 2021 Fundus CenterNet Diabetic Retinopathy Average Accuracy 98.10%
A132  [185] 2021 Ultrasound CenterNet Thyroid Nodule Accuracy, Mean IoU 77.92%, 79.58%
A133  [186] 2022 CT FoveaBox Lesion mAP 50.51%
A134  [187] 2021 MRI FoveaBox Lymph Nodes mAP 61.67%
A135  [188] 2022 CT CornerNet, CenterNet Blood vessels Precision, Recall 0.847, 0.865, 0.939, 0.703
A136  [144] 2021 Endoscopic CenterNet Gastric Polyps mAP, IoU 60.02%, 58.27%
A137  [189] 2023 MRI FoveaBox Lymphadenopathy mAP 52.3%
A138  [190] 2023 Radiographs Foveabox Lesion mAP 0.726
A139  [191] 2021 Endoscopic CornerNet, CenterNet Spine surgery Recall, Precision 0.796, 0.877, 0.839, 0.807
A140  [188] 2022 CT CornerNet, CenterNet Lower Extremities Precision 0.847, 0.939
A141  [192] 2021 Pahtological CornerNet, CenterNet Glomeruli AP 0.595, 0.574
A142  [177] 2023 Microscopic CenterNet Blood Cells F1, AP, mAP 0.854, 0.934, 0.788
A143  [193] 2023 CT CenterNet Rib fracture mAP 89.2%
A144  [194] 2020 Microscopic CenterNet Parasite eggs mAP, Weighted mAP 0.5000, 0.2612
A145  [195] 2022 Ultrasound CenterNet Clinical Diagnosis mAP 78.6%
A146  [196] 2021 Ultrasound CenterNet Breast cancer mAP 13.8%
A147  [172] 2022 Colonoscopy CenterNet Polyp detection Precision, Recall, F1 Score 95.27%, 93.35%, 94.30%
A148  [198] 2024 CT CenterNet Liver tumour Recall, Precision, mAP 0.8279, 0.8343, 0.8489
5.3.1 Analysis of Point-based Detectors
Similarly, the advanced detectors have been applied for the segmentation of different biomedical imaging modalities including MRI, X-Rays, dermoscopic images, colonoscopic images, microscopic images, OCTA, CT images, histopathological images, fundus images, ultrasound images, endoscopic images, and radiographs. If their performance is deeply analysed, the highest accuracies have been achieved using both cornernet and centernet with similar results in few cases and few fluctuations with each other. In some cases, cornernet outperforms centerent whereas in other cases, centernet outperforms cornernet. The results obtained from Foveabox are not very satisfactory in terms of accuracy. The major reasons behind this phenomenon of different results include size of datasets used in the experimentations, running environment, models parameters’ tuning, data divisions into training and testing sets, data augmentation approach, data pre-processing and post processing mechanisms among the most important factors.
6 CHALLENGES AND FUTURE DIRECTIONS
There are many challenges associated with different types of biomedical images as well the models applied for their segmentitons. This section outlines all the challenges suggest future directions related to these challenges and future trends in biomedical image segmentation used for diseases detection.
6.1 Challenges with Biomedical Image Segmentation
The segmentation of biomedical image segmentation face many challenges due to the complex nature of images data, variability of images, high power computational requirements, insufficient data availability, and different hardware and software requirements among other factors. The first and foremost challenge for the segmentation of biomedical images is the process of accruacte annotation and labeling. This is due to the fact that segmentation is not an easy task like classification which requires few values in the output rather it requires images in the output for preparing training datasets. This is impossible without the involvement of field expert along with high quality software tools which needs additional time, efforts, and cost. Secondly, the complex structure of biomedical images due to various shapes, formats, organs anatomy, cells structures and tissues structures have very deep effect on the performance of segmentation models. Another challenging task associated with the biomedical image segmentation is the scale of biomedical images which ranges from smaller nuclei of a cell to lareger organs of human body e.g., hearts, lungs, kidney, and other organs. Developing robust segmentation models for this large scale differences is highly challenging.
Another important and considerable issue related to biomedical images is the class imbalance problem  [198]. This issue arises due to unavailability of sufficient instances for all the available classes of diseases being considered in the problem. This problem has a very bad impact on the performance of the segmentation models which leads to higher inaccuracy in the segmentation process and biases in the prediction. Similarly, all types of biomedical images are affected by different types of noises which adequately degrade the quality of images. Other factors affecting the quality of medical images include low contrast, distortions, low quality illumination, and body movements during .capturing images. All these types of unwanted activities have negative impact on biomedical images processing resulting in poor performance of the segmentation tasks. To address all of these challenges, we need effective, efficient, and powerful communication and collaborations among different experts including medical experts, computer experts, and scientists. This will reduce the barriers prevailing in the processing of biomedical images which will improve the process of segmentation.
6.2 Challenges with Segmentation Models
There are different challenges associated with two stage detection models, one stage detection models, and the advanced detection models. All these challenges are highlighted in this section. The most important issue associated with two stage detection models is the complex architecture of the models due to performing the segmentation task in two stages namely region proposals and the classification. The second challenge with these models is their complete dependency on the mechanism of region proposal for generating bounding boxes which may result in the localization error which significantly decrease the segmentation accuracy of the models. Another main problem with the two stage detection models is the slow processing due to involvement of multiple stages for the segmentation task. Similarly, the models may be poorly trained due to optimization of two different components namely region proposals generation and the classification. To improve the convergence and ensuring the effectiveness in the learning both the training mechanism and the hyperparameter tuning process need to be highly careful and managed.
The major challenge of one stage detection model for biomedical image segmentation is low accuracy. The low accuracy results from the fact that these models directly perform the segmentation process with region proposals. Another major issue with one stage detectin model is their weakness in segmentation of small regions showing the abnormality.
Since, there is no specific mechanism for the identirication of different regions, therefore, it is very difficult for the detection to distinguish between the background and foreground of the biomedical images. Likewise, identification of different sizes, shapes, and textrure of various abnormal existence in the biomedical images is another challenge for one strage detection models. Another difficult mechanism for one stage detector is collecting contextual information from the surroundings of the affected areas which degrades the performance of the model in the segmentation process.
6.3 Future Research Directions
This review paper covers various object detection-based models for biomedical imaging segmentation. Although these models represent the latest architectures for biomedical image segmentation, the complexity of the segmentation process still leaves considerable room for improvement. This section outlines potential future research directions in biomedical imaging segmentation.
1) Multi-modal Approach: Current architectures for biomedical imaging segmentation typically rely on a uni-modal approach, where a model is used to segment images from a single modality. A significant direction for researchers is the application of a multi-modal approach, integrating information from multiple biomedical imaging modalities such as CT, X-Ray, MRI, and SPECT. Combining information from various modalities could greatly enhance the performance and accuracy of the segmentation process.
2) Transfer Learning: A major challenge in biomedical imaging segmentation is the scarcity of sufficient datasets, which complicates the segmentation process. This issue could be addressed through transfer learning, where models trained on large datasets are adapted for use on smaller datasets. This approach could provide a solid foundation for researchers and industries focused on biomedical imaging segmentation.
3) Advanced Deep Learning Models: While the deep learning models discussed in this review have contributed significantly to biomedical imaging segmentation, there is still potential for improvement. Advanced models, such as transformer-based architectures, Graph Neural Networks (GNN), hybrid models, and attention-based models, could further enhance segmentation accuracy, especially in images compromised by various types of noise.
4) Unsupervised and Semi-Supervised Approaches: Labeling and annotating biomedical images require significant resources, including time and effort. Unsupervised and semi-supervised learning approaches could alleviate some of these challenges. Models such as Self Organizing Maps (SOM), Generative Adversarial Networks (GANs), graph-based techniques, and autoencoders represent promising areas for future research.
5) Enhancement Techniques: Biomedical images often suffer from quality issues such as noise, poor illumination, and suboptimal image capture methods. Research into models and techniques that enhance image quality could improve segmentation accuracy and performance.
6) Clinical Aspects: Bridging the gap between technological advancements and clinical applicability is crucial. Future research should focus on developing methods that support the practical implementation of segmentation models, aiding in the detection of abnormalities across various human organs.
7) Hybrid approaches with large data: Exploring interdisciplinary approaches that combine biomedical imaging the data from other related domains such as bioinformatics, computational physics, and systems biology could also be explored to analyse the images based on different types of related data. For example, with the use of genetic data and integrating it in image segmentation could enhance the understanding of disease mechanisms at a molecular level, leading to more personalized and effective treatments.
8) Real-time Segmentation: Developing methods for real-time segmentation of biomedical images could significantly impact clinical diagnostics and surgical procedures. This would involve optimizing existing algorithms for speed without sacrificing accuracy, potentially utilizing edge computing technologies to process data directly on medical devices and providing instant results for clinical decision-making.
9) Interactive Segmentation: Enhancing segmentation results through user interaction is a promising avenue. Exploring the latest Large Language Models (LLMs) could provide new methods for user interaction with segmentation systems, potentially making them more intuitive and effective for various medical imaging tasks.
7 CONCLUSION
We conducted a comprehensive review of the deep learning based object detection methods use for biomedical image segmentation. Following, the standard SRL protocol, we selected 148 articles, and extracted information from these auricles. The extracted information include deep learning models, biomedical imaging modalities, and human disease. We have presented these information in tabular forms and provided various charts and classifications. Object detection based deep learning models are divided into three types namely two stage detectors, one-stage detectors, and advanced detectors which have been reviewed, critically analyzed. Finally, we highlighted different challenges associated with medical imaging modalities and object detection based deep learning models used for their segmentation, and highlighted future directions. To the best of our knowledge, this is the first SLR conducted on this specific topic within the field. We believe this work will serve as a foundational reference, providing essential guidance for new researchers and offering a comprehensive overview for both new and experienced researchers in the field.